# Whitehead’s problem and condensed mathematics

### v3

@@ Abstract.
One of the better-known independence results in general mathematics is Shelah’s solution to Whitehead’s problem of whether $\mathrm{Ext}^{1}(A,\mathbb{Z})=0$ implies that an abelian group $A$ is free. The point of departure for the present work is Clausen and Scholze’s proof that, in contrast, one natural interpretation of Whitehead’s problem within their recently-developed framework of condensed mathematics has an affirmative answer in $\mathsf{ZFC}$. We record two alternative proofs of this result, as well as several original variations on it, both for their intrinsic interest and as a springboard for a broader study of the relations between condensed mathematics and set theoretic forcing. We show more particularly how the condensation $\underline{X}$ of any locally compact Hausdorff space $X$ may be viewed as an organized presentation of the forcing names for the points of canonical interpretations of $X$ in all possible set-forcing extensions of the universe, and we argue our main result by way of this fact. We also show that when interpreted within the category of light condensed abelian groups, Whitehead’s problem is again independent of the $\mathsf{ZFC}$ axioms. In fact we show that it is consistent that Whitehead’s problem has a negative solution within the category of $\kappa$-condensed abelian groups for every uncountable cardinal $\kappa$, but that this scenario, in turn, is inconsistent with the existence of a strongly compact cardinal.
@@
The first author was supported by María Zambrano and Marie Skłodowska Curie (project 101110452: CatT) Fellowships at the University of Barcelona; the second author was supported by GAČR project 23-04683S and the Academy of Sciences of the Czech Republic (RVO 67985840); the third author was supported by GAČR project 23-05148S. A portion of this work was carried out while all three authors were participating in the Thematic Program on Set Theoretic Methods in Algebra, Dynamics and Geometry at the Fields Institute in spring of 2023. We thank the Fields Institute for their support and hospitality. The results of this paper were presented in detail at the Cornell Logic Seminar in Fall 2023. We would like to thank the participants of the seminar, particularly Mark Schachner and Justin Moore, for their engagement with the material and for a number of helpful remarks, and to thank Peter Scholze for valuable remarks on the paper as well.
@@ 1. Introduction
Clausen and Scholze’s proof of Theorem 1.2 is quite slick and relies on a structural analysis of the abelian subcategory of $\mathsf{CondAb}$ spanned by what are known as solid abelian groups. In this paper, we present a more elementary, combinatorial proof of Theorem 1.2 that is also more constructive in the sense that, given a nonfree abelian group $A$, we specify a highly natural associated Stone space $S$ and construct a nonzero element of $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)$. Our motivation for doing this is twofold. First, we seek to excavate some of the combinatorial and set theoretic ideas underlying Theorem 1.2 and thereby to improve our understanding of the mathematics surrounding Whitehead’s problem, in both general and condensed contexts. To this end, we prove the following theorem.
@@
    $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(2^{\kappa})\neq 0$;
  2. (2)
    in $V^{\mathbb{B}}$, $A$ is not Whitehead;
@@
Second, using Theorem 1.3, particularly clauses (2) and (3) thereof, as a starting point, we begin to investigate a more general connection between condensed mathematics and set theoretic forcing. We isolate a precise sense in which, given a locally compact Hausdorff space $X$, the condensed set $\underline{X}$ is simply an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions. We then exhibit a correspondence between the satisfaction of certain simple formulas concerning a topological space $X$ in two settings: first, in a Boolean-valued model $V^{\mathbb{B}}$, and second in the context of the space $C(S,X)$ of continuous functions from the Stone space $S$ of $\mathbb{B}$ to $X$. This correspondence, stated in Theorem 9.1, allows us, e.g., to immediately deduce clause (3) of Theorem 1.3 from clause (2) of the same theorem.
@@
    It is consistent with the axioms of $\mathsf{ZFC}$ that for every uncountable cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\kappa}}(\underline{A},\underline{\mathbb{Z}})=0$.
@@
Second, we extend Theorem 1.2 from the class of discrete abelian groups to the class of locally compact abelian groups and show that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})=0$ for a locally compact abelian group $A$ if and only if $A$ is projective in the category of locally compact abelian groups. We record as well several observations by Peter Scholze showing that Theorem 1.2 does not extend to the category of solid abelian groups.
The structure of the paper is as follows. In Section 2, we introduce the basic definitions and facts regarding condensed mathematics that we will need. In Section 3, we recall some classical algebraic facts pertinent to Whitehead’s problem and introduce their analogues in the condensed setting. In Section 4, we prove a slight strengthening of Stein’s aforementioned classical result [31] that all countable Whitehead groups are free. In Section 5, we prove clause (1) of Theorem 1.3, and, in Section 6, we adapt that proof to establish clause (2) of Theorem 1.3. From here we present two paths to clause (3) of Theorem 1.3. The first, recorded in Section 7, is by far the shorter and more direct of the two. The second, comprising Sections 8 and 9, passes through a more general investigation into the interplay between condensed mathematics and forcing. In Section 8, a robust correspondence is established between
@@
Section 9 then establishes the aforementioned equivalence between the settings of $C(S,X)$ and $V^{\mathbb{B}}$ with respect to interpretations of certain formulas. We expect the results of this section to be applicable in a wide variety of contexts, and derive clause (3) of Theorem 1.3 as an immediate corollary. In Section 10, we prove Theorem 1.4(1). In Section 11, we record an algebraic lemma converting the groups $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)$ into a more tractable, and suggestive, form. This lemma yields an alternative proof of clauses (1) and (3) of Theorem 1.3 (and hence of Theorem 1.2), as well as clauses (2) and (3) of Theorem 1.4. In Section 12 we prove an extension of Theorem 1.2 to the class of locally compact abelian groups, and record impediments to its extension to the class of solid abelian groups. We close out our work in Section 13 with a few concluding remarks and one main open question.
@@ 1.1. Notation and conventions
Given a Boolean algebra $\mathbb{B}$, we let $S(\mathbb{B})$ denote the Stone space of $\mathbb{B}$. Concretely, $S(\mathbb{B})$ is the space of ultrafilters on $\mathbb{B}$, with basic clopen sets given by
$N_{b}:=\{U\in S(\mathbb{B})\mid b\in U\}$
for $b\in\mathbb{B}$.
@@ 2. Condensed abelian groups
    is a bijection.
@@
    is a bijection.
@@
Given a Hausdorff space $X$, one can define a condensed set $\underline{X}$ by letting $\underline{X}(S)=C(S,X)$, i.e., $X(S)$ is the set of continuous functions from $S$ to $X$ for all $S\in\mathsf{Prof}$; we will sometimes term this $\underline{X}$ the condensation of $X$. This operation defines a functor from the category $\mathsf{Haus}$ of Hausdorff topological spaces to $\mathsf{Cond}$. When restricted to the full subcategory of $\mathsf{Haus}$ spanned by the compactly generated Hausdorff spaces, it is a fully faithful embedding (cf. [26, Proposition 1.7]).
@@
For each uncountable cardinal $\kappa$, there is an obvious forgetful functor from $\mathsf{CondAb}_{\kappa}$ to $\mathsf{Cond}_{\kappa}$, and this functor admits a left adjoint that sends a $\kappa$-condensed set $T$ to a $\kappa$-condensed abelian group denoted $\mathbb{Z}[T]$. Concretely, as noted in [26, §2], $\mathbb{Z}[T]$ is the sheafification of the functor from $\mathsf{ED}_{\kappa}$ to $\mathsf{Ab}$ that sends each $S\in\mathsf{ED}_{\kappa}$ to the free group on $T(S)$. Moreover, $\{\mathbb{Z}[\underline{S}]\mid S\in\mathsf{ED}_{\kappa}\}$ is, for any strong limit cardinal $\kappa$, a set of compact projective generators for $\mathsf{CondAb}_{\kappa}$. These facts straightforwardly propagate upward to the category $\mathsf{CondAb}$ of all condensed abelian groups.
@@ 3. Algebraic preliminaries
[…]
in which both $F$ and $K$ are free abelian groups. In such resolutions, we will often implicitly assume that $K$ is a subgroup of $F$ and $\iota$ is the inclusion map. We note that every infinite abelian group $A$ has a canonical free resolution, in which $F$ is the free abelian group generated by the underlying set of $A$ and $K$ is the kernel of the induced surjection from $F$ onto $A$. The elements of $K$ are therefore the formal linear combinations of elements of $F$ that evaluate to 00 in $A$.
@@
Proposition 3.1 (Clausen–Scholze, [26, Proposition 4.2]).
@@
As noted in [26, proof of Corollary 4.8] (and as follows from items (ii) and (iii) of its page 13),
@@
for any condensed abelian groups $T,M$ and any extremally disconnected profinite set $S$. When $S=\ast$, our general claim follows from the facts that $\mathbb{Z}[*]=\underline{\mathbb{Z}}$ and that $\underline{\mathbb{Z}}$ is the unit object for the condensed tensor product $\otimes$. To see the assertion for discrete abelian groups, recall that their category fully and faithfully embeds into $\mathsf{CondAb}$, and observe by [18, Props. 2.18 and 2.19] that if $\cdots\to P^{1}\to P^{0}\to A$ is a projective (hence free) resolution of $A$ then $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to\underline{A}$ is a projective resolution of $\underline{A}$. Writing $\mathcal{P}$ and $\underline{\mathcal{P}}$ for $\cdots\to P^{1}\to P^{0}\to 0$ and $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to 0$, respectively, we then have
@@
It is natural at this point in the discussion to declare a condensed abelian group $T$ to be Whitehead if $\underline{\mathrm{Ext}}^{1}(T,\underline{\mathbb{Z}})=0$. Observe next that, in close analogy with the classical situation, if $F$ is a free (discrete) abelian group, then $\underline{F}$ is Whitehead in $\mathsf{CondAb}$. To see this, let $\kappa$ be a cardinal, and let $F=\bigoplus_{\kappa}\mathbb{Z}$ be the free group on $\kappa$ generators; it follows then that $\underline{F}\cong\bigoplus_{\kappa}\underline{\mathbb{Z}}$, where the direct sum on the right is computed in $\mathsf{CondAb}$, from [18, Proposition 2.19]. By [18, Remark 3.10 and Lemma 4.2], we then have
@@
where ¯$\underline{\mathbb{Z}}[0]$ denotes the chain complex that is $\underline{\mathbb{Z}}$ in degree 00 and 00 elsewhere. In particular, we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{F},\underline{\mathbb{Z}})=0$.
@@
In the coming sections, we will give a more concrete, combinatorial proof of Theorem 1.2. In particular, given a nonfree discrete abelian group $A$, we will identify a natural profinite set $S$ such that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\neq 0$. The process by which this will be done is analogous to that described in the classical setting at the beginning of this section, so in the interest of symmetry, we end this section by giving an outline of the structure of the argument.
Suppose that $A$ is a nonfree discrete abelian group, and let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be a free resolution of $A$. Applying $\underline{\mathrm{Hom}}(\cdot,\underline{\mathbb{Z}})$ to the image of this short exact sequence in $\mathsf{CondAb}$ yields the exact sequence
$0\rightarrow\underline{\mathrm{Hom}}(\underline{A},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Hom}}(\underline{F},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Hom}}(\underline{K},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})\rightarrow 0$
Now, evaluating this sequence at a profinite set $S$ and recalling Proposition 3.1, we obtain
$0\rightarrow\underline{\mathrm{Hom}(A,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(K,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\rightarrow 0$
@@
Therefore, to show that $\underline{A}$ is not Whitehead, it will suffice to find a profinite set $S$ and a continuous $\varphi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ that does not “lift pointwise” to a map from $S$ to $\mathrm{Hom}(F,\mathbb{Z})$ (i.e., for which there is no continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ for which $\psi(s)\restriction K=\varphi(s)$ for all $s\in S$). The construction of such an $S$ and $\varphi$ will be the subject of the next sections.

### v5

@@ Abstract.
One of the better-known independence results in general mathematics is Shelah’s solution to Whitehead’s problem of whether $\mathrm{Ext}^{1}(A,\mathbb{Z})=0$ implies that an abelian group $A$ is free. The point of departure for the present work is Clausen and Scholze’s proof that, in contrast, one natural interpretation of Whitehead’s problem within their recently-developed framework of condensed mathematics has an affirmative answer in $\sf ZFC$. We record two alternative proofs of this result, as well as several original variations on it, both for their intrinsic interest and as a springboard for a broader study of the relations between condensed mathematics and set theoretic forcing. We show more particularly how the condensation $\underline{X}$ of any locally compact Hausdorff space $X$ may be viewed as an organized presentation of the forcing names for the points of canonical interpretations of $X$ in all possible set-forcing extensions of the universe, and we argue our main result by way of this fact. We also show that when interpreted within the category of light condensed abelian groups, Whitehead’s problem is again independent of the $\sf ZFC$ axioms. In fact we show that it is consistent that Whitehead’s problem has a negative solution within the category of $\kappa$-condensed abelian groups for every uncountable cardinal $\kappa$, but that this scenario, in turn, is inconsistent with the existence of a strongly compact cardinal.
@@
The first author was supported by María Zambrano and Marie Skłodowska Curie (project 101110452: CatT) Fellowships at the University of Barcelona; the second author was supported by GAČR project 23-04683S and the Academy of Sciences of the Czech Republic (RVO 67985840); the third author was supported by GAČR project 23-05148S
@@ 1. Introduction
Clausen and Scholze’s proof of Theorem 1.2 is quite slick and relies on a structural analysis of the abelian subcategory of $\mathsf{CondAb}$ spanned by what are known as solid abelian groups. In this paper, we present a more elementary, combinatorial proof of Theorem 1.2 that is also more constructive in the sense that, given a nonfree abelian group $A$, we specify a highly natural associated Stone space $S$ and construct a nonzero element of $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)$. Our motivation for doing this is twofold. First, we seek to excavate some of the combinatorial and set theoretic ideas underlying Theorem 1.2 and thereby to improve our understanding of the mathematics surrounding Whitehead’s problem, in both classical and condensed contexts. To this end, we prove the following theorem, item (2) of which strengthens Theorem 1.2.
@@
    in $V^{\mathbb{B}}$, $A$ is not Whitehead; and
@@
Second, using Theorem 1.3 as a starting point, we begin to investigate a more general connection between condensed mathematics and set theoretic forcing. We isolate a precise sense in which, given a locally compact Hausdorff space $X$, the condensed set $\underline{X}$ is simply an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions. We then exhibit a correspondence between the satisfaction of certain simple formulas concerning a topological space $X$ in two settings: first, in a Boolean-valued model $V^{\mathbb{B}}$, and second in the context of the space $C(S,X)$ of continuous functions from the Stone space $S$ of $\mathbb{B}$ to $X$. This correspondence, stated in Theorem 9.1, allows us, e.g., to immediately deduce clause (2) of Theorem 1.3 from clause (1) of the same theorem.
@@
    It is consistent with the axioms of $\sf ZFC$ that for every uncountable cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\kappa}}(\underline{A},\underline{\mathbb{Z}})=0$.
@@
Second, we extend Theorem 1.2 from the category of (discrete) abelian groups to the category of locally compact abelian groups; here it is useful to recall that the projective objects of the former category are exactly the free ones. In particular, we prove the following theorem:
Theorem 1.5.
if and only if $A$ is projective in $\mathsf{LCA}$.
We record as well several observations by Peter Scholze showing that Theorem 1.2 does not extend to the category of solid abelian groups.
The structure of the paper is as follows. In Section 2, we introduce the basic definitions and facts regarding condensed mathematics that we will need. In Section 3, we recall some classical algebraic facts pertinent to Whitehead’s problem and introduce their analogues in the condensed setting. In Section 4, we prove a slight strengthening of Stein’s aforementioned classical result [33] that all countable Whitehead groups are free. In Section 5, we prove a precursor to Theorem 1.3 concerning the space $2^{\kappa}$, and, in Section 6, we adapt that proof to establish clause (1) of Theorem 1.3. From here we present two paths to clause (2) of Theorem 1.3. The first, recorded in Section 7, is by far the shorter and more direct of the two. The second, comprising Sections 8 and 9, passes through a more general investigation into the interplay between condensed mathematics and forcing. In Section 8, a robust correspondence is established between
@@
Section 9 then establishes the aforementioned equivalence between the settings of $C(S,X)$ and $V^{\mathbb{B}}$ with respect to interpretations of certain formulas. We expect the results of this section to be applicable in a wide variety of contexts, and derive clause (2) of Theorem 1.3 as an immediate corollary. In Section 10, we prove Theorem 1.4(1). In Section 11, we record an algebraic lemma converting the groups $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)$ into a more tractable, and suggestive, form. This lemma yields an alternative proof of clause (2) of Theorem 1.3 (and hence of Theorem 1.2), as well as clauses (2) and (3) of Theorem 1.4. In Section 12 we prove Theorem 1.5 and record impediments to the extension of Theorem 1.2 to the class of solid abelian groups. We close out our work in Section 13 with a few concluding remarks and one main open question.
@@ 1.1. Notation and conventions
Given a Boolean algebra $\mathbb{B}$, we let $\mathbb{B}^{+}$ denote the set of all nonzero elements of $\mathbb{B}$ and $S(\mathbb{B})$ denote the Stone space of $\mathbb{B}$. Concretely, $S(\mathbb{B})$ is the space of ultrafilters on $\mathbb{B}$, with basic clopen sets given by
$N_{b}:=\{\mathcal{U}\in S(\mathbb{B})\mid b\in\mathcal{U}\}$
for $b\in\mathbb{B}^{+}$.
@@ 2. Condensed abelian groups
    induced by the inclusion maps of $S_{0}$ and $S_{1}$ into $S_{0}\sqcup S_{1}$ is a bijection.
@@
    induced by $f$ is a bijection. [^2] In slightly more detail, $f$ induces a map from $T(S)$ to $T(S^{\prime})$ whose range is readily seen to be contained in the set $\{x\in T(S^{\prime})\mid p_{0}^{*}(x)=p_{1}^{*}(x)\}$. This condition then asserts that this map is a bijection between $T(S)$ and $\{x\in T(S^{\prime})\mid p_{0}^{*}(x)=p_{1}^{*}(x)\}$.
@@
Given a Hausdorff space $X$, one can define a condensed set $\underline{X}$ by letting $\underline{X}(S)=C(S,X)$, i.e., $X(S)$ is the set of continuous functions from $S$ to $X$ for all $S\in\mathsf{Prof}$; we will sometimes term this $\underline{X}$ the condensation of $X$. This operation defines a functor from the category $\mathsf{Haus}$ of Hausdorff topological spaces to $\mathsf{Cond}$. When restricted to the full subcategory of $\mathsf{Haus}$ spanned by the compactly generated Hausdorff spaces, it is a fully faithful embedding (cf. [28, Proposition 1.7]).
@@
For each uncountable cardinal $\kappa$, there is an obvious forgetful functor from $\mathsf{CondAb}_{\kappa}$ to $\mathsf{Cond}_{\kappa}$, and this functor admits a left adjoint that sends a $\kappa$-condensed set $T$ to a $\kappa$-condensed abelian group denoted $\mathbb{Z}[T]$. Concretely, as noted in [28, §2], $\mathbb{Z}[T]$ is the sheafification of the functor from $\mathsf{Prof}_{\kappa}$ to $\mathsf{Ab}$ that sends each $S\in\mathsf{Prof}_{\kappa}$ to the free group on $T(S)$. Moreover, $\{\mathbb{Z}[\underline{S}]\mid S\in\mathsf{ED}_{\kappa}\}$ is, for any strong limit cardinal $\kappa$, a set of compact projective generators for $\mathsf{CondAb}_{\kappa}$. These facts straightforwardly propagate upward to the category $\mathsf{CondAb}$ of all condensed abelian groups.
@@ 3. Algebraic preliminaries
$0\rightarrow K\xrightarrow{\iota}F\xrightarrow{\pi}A\rightarrow 0$
in which both $F$ and $K$ are free abelian groups. In such resolutions, we will often implicitly assume that $K$ is a subgroup of $F$ and $\iota$ is the inclusion map. We note that every infinite abelian group $A$ has a canonical free resolution, in which $F$ is the free abelian group generated by the underlying set of $A$ and $K$ is the kernel of the induced surjection from $F$ onto $A$. The elements of $K$ are therefore the formal linear combinations of elements of $F$ that evaluate to 0 in $A$.
@@
Proposition 3.1 (Clausen–Scholze, [28, Proposition 4.2]).
@@
As noted in [28, proof of Corollary 4.8] (and as follows from items (ii) and (iii) of its page 13),
@@
for any condensed abelian groups $T,M$ and any extremally disconnected profinite set $S$. When $S=\ast$, our general claim follows from the facts that $\mathbb{Z}[*]=\underline{\mathbb{Z}}$ and that $\underline{\mathbb{Z}}$ is the unit object for the condensed tensor product $\otimes$. To see the assertion for discrete abelian groups, recall that their category fully and faithfully embeds into $\mathsf{CondAb}$, and observe by [19, Props. 2.18 and 2.19] that if $\cdots\to P^{1}\to P^{0}\to A$ is a projective (hence free) resolution of $A$ then $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to\underline{A}$ is a projective resolution of $\underline{A}$. Writing $\mathcal{P}$ and $\underline{\mathcal{P}}$ for $\cdots\to P^{1}\to P^{0}\to 0$ and $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to 0$, respectively, we then have
@@
It is natural at this point in the discussion to declare a condensed abelian group $T$ to be Whitehead if $\underline{\mathrm{Ext}}^{1}(T,\underline{\mathbb{Z}})=0$. Observe next that, in close analogy with the classical situation, if $F$ is a free (discrete) abelian group, then $\underline{F}$ is Whitehead in $\mathsf{CondAb}$. To see this, let $\kappa$ be a cardinal, and let $F=\bigoplus_{\kappa}\mathbb{Z}$ be the free group on $\kappa$ generators; it follows from [19, Proposition 2.19] that $\underline{F}\cong\bigoplus_{\kappa}\underline{\mathbb{Z}}$, where the direct sum on the right is computed in $\mathsf{CondAb}$. By [19, Remark 3.10 and Lemma 4.2], we then have
@@
where $\underline{\mathbb{Z}}[0]$ denotes the chain complex that is $\underline{\mathbb{Z}}$ in degree 0 and 0 elsewhere. In particular, we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{F},\underline{\mathbb{Z}})=0$.
@@
In the coming sections, we will give a more concrete, combinatorial proof of Theorem 1.2. In particular, given a nonfree discrete abelian group $A$, we will identify a natural $S\in\mathsf{ED}$ such that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\neq 0$. The process by which this will be done is analogous to that described in the classical setting at the beginning of this section, so in the interest of symmetry, we end this section by giving an outline of the structure of the argument.
Suppose that $A$ is a nonfree discrete abelian group, and let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be a free resolution of $A$. Using the discreteness of the groups under consideration, it is straightforward to observe that the induced sequence $0\rightarrow\underline{K}\rightarrow\underline{F}\rightarrow\underline{A}\rightarrow 0$ is exact in $\mathsf{CondAb}$. Applying $\underline{\mathrm{Hom}}(\cdot,\underline{\mathbb{Z}})$ to this short exact sequence yields the exact sequence
$0\rightarrow\underline{\mathrm{Hom}}(\underline{A},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Hom}}(\underline{F},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Hom}}(\underline{K},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})\rightarrow 0.$
Given an $S\in\mathsf{ED}$, the functor from $\mathsf{CondAb}$ to $\mathsf{Ab}$ sending $T\in\mathsf{CondAb}$ to $T(S)\in\mathsf{Ab}$ is exact (cf. [16, Lemma 6.1.3]). Thus, evaluating the above sequence at a space $S\in\mathsf{ED}$ and recalling Proposition 3.1, we obtain
$0\rightarrow\underline{\mathrm{Hom}(A,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(K,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\rightarrow 0.$
@@
Therefore, to show that $\underline{A}$ is not Whitehead, it will suffice to find a space $S\in\mathsf{ED}$ and a continuous $\varphi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ that does not “lift pointwise” to a map from $S$ to $\mathrm{Hom}(F,\mathbb{Z})$, i.e., for which there is no continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ for which $\psi(s)\restriction K=\varphi(s)$ for all $s\in S$. The construction of such an $S$ and $\varphi$ will be the subject of the next sections. For expository reasons, and because it more closely derives from classical arguments around Whitehead’s problem, we begin by finding $S$ and $\varphi$ as above but for which $S$ is merely profinite and not extremally disconnected. We then show how to improve these arguments to obtain $S\in\mathsf{ED}$, as desired.